The Green Giant has a 7 percent profit margin and a 38 percent divided payout ratio. the total asset turnover is 1.3 times and the equity multiplier is 1.4 times. what is the sustainable rate if growth?
Answer #1
Return on equity = Profit margin*Total assets turnover* Equity multiplier
7% *1.3*1.4 = 12.74%
Retention ratio = 1- Payout ratio = 1 - 0.38 = 62%
Sustainable growth rate = Return on equity * Retention ratio
12.74*0.62 = 7.90%